Need this asap please. thanks

Sun Microsystems, is a manufacturer of mid-size computers, which are primarily used by medium and small businesses. Sun Microsystems’s product line currently consists of three models of mid-size computers. The following data are available regarding the models:

Model

Selling Price per Unit

Variable Cost per Unit

Demand/Year (units)

Model SMX1

$1,500

$1,000

2,000

Model SMX2

$2,000

$1,300

1,000

Model SMX3

$2,500

$1,400

500

Sun Microsystems is considering the addition of a fourth model to its line of mid-size computers.

This model, the SMX-4 would be sold for $3,000.  
The variable cost of this unit is the $101
The demand for the new Model SMX4 is estimated to be 600 units per year.  
Fifty percent (50%) of the new model’s unit sales are expected to come from other models already being manufactured by Sun Microsystems
60% of the cannibalized amount from Model SMX-1,
   10% from Model SMX-2,
   30% from Model SMX-3.
Sun Microsystems will incur a fixed cost of $500,000 to add the new model SMX-4 to its product line.

The president of Sun Microsystems will only fund projects which show a positive cash flow by the end of the first year of the project.   

Answer the following questions:

  1. What is unit contribution for SMX-4?
  1. Based on the above data should Sun Microsystems add the new Model SMX-4 to its line of mid-size computers?

  1. Why (Support answer with increase/decrease in total contribution)?
Answer #1
Unit Contribution margin for SMX 4 =$3,000 - $101 =$2,899 per unit
Contribution margin of existing model:
SMX 1 =$1,500 - $1,000 =$500
SMX 2 =$2,000 - $1,300 =$700
SMX 3 =$2,500 - $1,400 =$1,100
Additional Contribution margin from sale of SMX 4 =600*$2,899 =$1,739,400(A)
Total unit sale loss of existing model =600*50% =300 units
Lost Contribution margin of SMX 1 =300*60%*$500 =$90,000
Lost Contribution margin of SMX 2 =300*10%*$700 =$21,000
Lost Contribution margin of SMX 3 =300*30%*$1,100 =$99,000
Total Contribution margin lost =$210,000(B)
Additional Fixed cost to add new model =$500,000©
Net Cash flowby end of 1st year of project =$1,739,400 - $210,000 - $500,000 =$1,029,400
Since the Net cash flow is positive hence the SMX 4 should be introduced